OpenStudy (anonymous):

write the equation of the line that passes through (–1, 4) with a slope of –1 in standard form.

OpenStudy (anonymous):

helppp pleaseeeee anyone?

OpenStudy (anonymous):

which way did you learn to do, slope form or point slope?

OpenStudy (anonymous):

i honestly have no idea. thats why i need help lol

OpenStudy (anonymous):

ok so y=mx+b or y-y1=m(x-x1)

OpenStudy (anonymous):

y=mx+b i thinkk

OpenStudy (anonymous):

ok so if m= the slope you could plug in the slope and the point in the equation (x,y) and solve for b

OpenStudy (anonymous):

do you know what the answer would be?

OpenStudy (anonymous):

if you do the proper steps you should know. Are you not understanding?

OpenStudy (anonymous):

nah not really. im doing a class online and i need the answeer for a test

OpenStudy (anonymous):

you shouldnt be taking a test on something you dont know but ok, give me a sec to do it

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

y+x= 2 i thinkkk

OpenStudy (anonymous):

c

OpenStudy (anonymous):

np

OpenStudy (anonymous):

Write the equation of the line that passes through (1, 3) and (4, 4) in standard form. what about this question/?